OpenStudy (anonymous):

PLEASE help im not sure how to solve this.... Computer Craft pays its sales associates a 15% commission on every sale. Barry sold a laptop for $1,900 and a printer for $300 How much money did Barry earn on the sale of the items?

OpenStudy (anonymous):

HEEELLLLPPPPP??? im confused on the process for solving

jimthompson5910 (jim_thompson5910):

The total comes to 1900+300 = 2200 dollars

OpenStudy (anonymous):

correct then do i multiply that by .15?

jimthompson5910 (jim_thompson5910):

Now take 15% of this 15% of 2200 = (15/100)*2200 15% of 2200 = 0.15*2200 15% of 2200 = 330 So he earned $330

jimthompson5910 (jim_thompson5910):

yes, you got it, you multiply the total by 0.15

OpenStudy (anonymous):

(6x^3+18x^2)-3x^2=???

jimthompson5910 (jim_thompson5910):

(6x^3+18x^2)-3x^2 6x^3+18x^2-3x^2 6x^3+(18x^2-3x^2) Keep going

OpenStudy (anonymous):

6x^3+(15x^2)

jimthompson5910 (jim_thompson5910):

or 6x^3 + 15x^2

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

can you help me again im trying to solve this by myself haha

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

6(4^2-5)-5^2

OpenStudy (anonymous):

so do i subtract the 5 in parenthesis first or the similar ones first then do that?

jimthompson5910 (jim_thompson5910):

you square each term first

jimthompson5910 (jim_thompson5910):

then you do the subtraction in the parenthesis then you multiply then you finally subtract

OpenStudy (anonymous):

i got 8

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

6(4^2-5)-5^2 6(16-5)-25 6(11)-25 66-25 41 So the answer is actually 41

OpenStudy (anonymous):

oh i just did 4 times 2 instead of 4 times 4

jimthompson5910 (jim_thompson5910):

yes, 4^2 = 4*4 = 16

OpenStudy (anonymous):

i have another question for you

jimthompson5910 (jim_thompson5910):

ask away

OpenStudy (anonymous):

6x divided by 3y minus 4 if x is 20 and y is 2 but i am confused because then it would be 620 divided by 32 right or is that wrong

jimthompson5910 (jim_thompson5910):

\[\Large \frac{6x}{3y-4}\] OR \[\Large \frac{6x}{3y}-4\] ??

OpenStudy (anonymous):